OpenStudy (anonymous):

write an equation for the line that is parallel to the given line and passes through the given point y=x+8;(2,16)? A. y=x B. y=x+6 C. y=x-6 ***I got B. but I am not sure. Need help!!!

OpenStudy (texaschic101):

Are you sure you wrote the right answer choices down ? y = mx + b -- the slope is going to be the same, it will be 1 y = 1x + b using points (2,16) 16 = 1(2) + b 16 = 2 + b 16 - 2 = b 14 = b The parallel line would be : y = x + 14